In the laboratory a student finds that it takes 145 Joules to increase the temperature of 10.6 grams of gaseous neon from 21.2 to 35.4 degrees Celsius.

The specific heat of neon calculated from her data is
J/g°C.
1
Expert's answer
2020-12-01T09:34:24-0500

Q=mc∆T.

c= Q/m∆T = 145/10.6 x 14.2 = 0,963327132 J/g°C.